### v1

# Reasoning by Exploration: A Unified Approach to Retrieval and Generation over Graphs

Abstract.
Reasoning over structured graphs remains a fundamental challenge for Large Language Models (LLMs), particularly when scaling to large graphs. Existing approaches typically follow the retrieval-augmented generation (RAG) paradigm: first retrieving subgraphs relevant to the query and then generating answers conditioned on the retrieved subgraphs. However, such two-phase pipelines often struggle to faithfully incorporate graph structure, since the generation process is ultimately constrained by the quality and completeness of the retrieved subgraph. Although many advanced retrievers have been proposed recently to mitigate this issue, they are usually tailored to the training graphs and generalize poorly to unseen graphs, which limits their practical applicability. In this work, we propose Reasoning by Exploration (RoE), a novel approach that unifies retrieval and generation by framing reasoning over graphs as a process of graph exploration. At each step, the LLM selects candidate nodes and edges to explore, gradually constructing reasoning paths and generating answers along the way. To enable effective exploration, RoE is trained in two stages: supervised fine-tuning (SFT) on gold reasoning paths, followed by reinforcement learning (RL) to enhance exploration effectiveness and generalization. Experiments on benchmark datasets demonstrate that RoE achieves substantial overall improvements over baselines, while also generalizing effectively to unseen graphs.
Large Language Models, Graph Reasoning, Retrieval-Augmented Generation, Multi-hop Question Answering
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NY
1. Introduction
Large language models (LLMs) (Chang et al., 2024; Thirunavukarasu et al., 2023; Kasneci et al., 2023) have demonstrated impressive abilities across a wide range of natural language processing tasks. To further enhance their reliability and factuality, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Gao et al., 2023; Jiang et al., 2023) has become a widely adopted paradigm. By retrieving external knowledge before generating responses, RAG has achieved strong performance in diverse domains, such as healthcare (Xu et al., 2024), law (Wiratunga et al., 2024), finance (Zhang et al., 2023), and education (Chu et al., 2025). However, most existing RAG systems operate over unstructured text, retrieving passages through lexical or semantic similarity, and therefore do not fully leverage the relational structure inherent in many forms of knowledge.
Graphs provide a natural representation for structured knowledge, capturing rich relational dependencies among entities. They are widely used in domains such as knowledge bases, social networks, and scientific applications including biology and chemistry (Wu et al., 2020; Ma and Tang, 2021; Wu et al., 2023a). Recently, Graph Retrieval-Augmented Generation (GraphRAG) has emerged as a promising way to integrate such graph-structured knowledge into LLMs (Han et al., 2024; Peng et al., 2024). There are mainly two steps in GraphRAG: first, a retriever selects query-relevant content from the entire graph, where the retrieved units may be nodes, paths, or subgraphs; then, a generator, usually an LLM, produces answers conditioned on the retrieved subgraphs. This framework has been successfully applied to tasks such as knowledge graph question answering (KGQA) (Yasunaga et al., 2021, 2022; Zhang et al., 2022a), where answering a query often requires multi-step reasoning over the graphs.
While efficient, this two-phase pipeline may constrain reasoning quality due to the disjoint optimization of retriever and generator. First, the retriever often ignores the specific needs of the generator, which can lead to missing critical evidence or introducing irrelevant noise. For example, consider the multi-hop question “Who are the nephews of Winston Churchill?”. This query may have multiple answers and can be reached through multiple reasoning paths, such as “Churchill $\rightarrow$ sibling $\rightarrow$ child” or “Churchill $\rightarrow$ parent $\rightarrow$ child $\rightarrow$ child”. Heuristic-based retrievers (Yasunaga et al., 2022; Feng et al., 2020) (e.g., retrieving fixed $k$-hop subgraphs) or GNN-based retrievers (Mavromatis and Karypis, 2024; Liu et al., 2024) may include a large amount of irrelevant information about Churchill, overwhelming the generator with noise (Guo et al., 2025a). In contrast, LLM-based retrievers (Luo et al., 2023; Wu et al., 2023b) may generate only a limited set of reasoning paths, leading to missing valid entities and incomplete answers, as shown in Section 3. Second, retrievers or generators trained on specific graphs or datasets usually generalize poorly to unseen graphs as shown in Section 3. For instance, a path-based retriever tuned on a movie knowledge base may learn to generate reasoning paths primarily around relations such as “actor $\rightarrow$ film $\rightarrow$ director”, but it struggles to generalize when applied to graphs with very different relational structures, such as biomedical graphs involving “gene $\rightarrow$ protein $\rightarrow$ disease” paths. This dependence leads to sharp performance drops when deployed on unseen graphs, limiting the robustness and real-world applicability of current GraphRAGs.
To address these limitations, we propose Reasoning by Exploration (RoE), a unified framework that leverages the reasoning ability of LLMs to perform retrieval and generation simultaneously within the same model. Instead of separating the two stages, RoE frames reasoning as a step-by-step graph exploration process, where the model incrementally expands nodes and edges while constructing reasoning paths and generating answers in a unified manner. In this way, RoE has the access to the whole graph. This design directly mitigates the problem of missing or noisy retrieval, since the model actively decides which parts of the whole graph to explore based on its evolving reasoning process, instead of relying on a static retriever. However, it faces two main challenges. First, how can we equip LLMs, which pretrained primarily on unstructured text, with the ability to perform structured graph exploration? Second, how can we ensure that the model learns generalizable exploration strategies rather than memorizing specific graph patterns, so that it can adapt to unseen graphs? To tackle these challenges, we adopt a two-stage training strategy. In the first stage, we construct gold exploration trajectories and apply supervised fine-tuning (SFT) to teach the LLM to expand nodes and edges step by step. In the second stage, we employ reinforcement learning (RL) with different reward signals to further refine the exploration policy, encouraging the model to efficiently discover valid reasoning paths while improving its ability to generalize across diverse and unseen graphs. This combination enables RoE to leverage both explicit supervision and adaptive learning signals, resulting in more faithful reasoning and stronger generalization to unseen graphs. Our key contributions are summarized as follows:
  • We introduce Reasoning by Exploration (RoE), a unified framework that integrates retrieval and generation into a single process by framing reasoning as graph exploration, allowing LLMs to construct reasoning paths and answers simultaneously.
  • We develop a two-stage training strategy that equips LLMs with exploration ability: supervised fine-tuning (SFT) on gold reasoning trajectories to learn step-wise expansion, followed by reinforcement learning (RL) to refine exploration policies and enhance generalization to unseen graphs.
  • We conduct extensive experiments on multi-hop reasoning benchmarks, demonstrating that RoE significantly outperforms state-of-the-art baselines while exhibiting strong generalization across diverse unseen graphs.
2. Related Works
2.1. GraphRAG
GraphRAG (Han et al., 2024; Peng et al., 2024) aims to retrieve relevant information from external graphs to improve the performance and reduce the hallucination of LLMs. Compared to traditional RAG systems (Lewis et al., 2020; Gao et al., 2023) that operate on unstructured text, GraphRAG leverages the relational structure of graphs, where nodes represent entities and edges encode relations between nodes. Similar to RAG, GraphRAG generally consists of two main components: a retriever, which selects query-relevant graph content, and a generator, typically an LLM, which produces answers based on the retrieved subgraphs. To accommodate the graph structure, there are mainly three types of retrievers:
Heuristic-based Retriever. Heuristic-based retrievers typically begin by using similarity-based methods (Wen et al., 2023; Sanmartin, 2024) or entity and relation extraction techniques (Al-Moslmi et al., 2020; Feng et al., 2020) to identify seed nodes and edges that are relevant to the query. These seeds are then expanded by retrieving their $k$-hop neighbors (Yasunaga et al., 2021; Choudhary and Reddy, 2023), forming a candidate subgraph for reasoning. A representative example is G-Retriever (He et al., 2024), which first leverages embedding models to retrieve the most similar nodes and edges to the query, and then applies the Prize-Collecting Steiner Tree (PCST) algorithm to extract a compact subgraph centered around the selected seeds.
GNN-based Retriever. GNN-based retrievers (Liu et al., 2024; Fang et al., 2024) typically train Graph Neural Networks (GNNs) (Wu et al., 2020; Ma and Tang, 2021) to model retrieval as a node classification task, where the correct answer entities are labeled as positive and others as negative. For example, GNN-RAG (Mavromatis and Karypis, 2024) first trains a GNN model and then retrieves the nodes with the highest probability scores as candidate answers, while also identifying the shortest paths connecting the question entities with these candidates to construct supporting evidence. Compared to heuristic-based methods, GNN-based retrievers can capture richer relational dependencies and are more adaptive to complex queries.
LLM-based Retriever. LLM-based retrievers leverage the reasoning ability of LLMs to generate retrieval plans instead of relying solely on graph embeddings or heuristics. For example, they may prompt an LLM to produce SQL queries over structured knowledge bases (Jiang et al., 2024) or to explicitly plan multi-hop paths from the query entities to answers (Zhu et al., 2024). A representative method is RoG (Luo et al., 2023), which fine-tunes an LLM to generate reasoning relation paths and then retrieves the corresponding paths from the graph.
While these retrievers have demonstrated success across different tasks, they remain unaware of the generation process, which often leads to incomplete retrieval and the introduction of substantial noise as shown in Section 3. Moreover, training-based methods are usually tailored to a specific graph, further limiting their ability to generalize to unseen graphs. Instead, the proposed RoE unifies retrieval and generation within the same process, enabling the model to actively explore the graph while constructing reasoning paths and generating answers in an integrated manner.
2.2. Reasoning of Large Language Models
Large Language Models (LLMs) have demonstrated strong performance on natural language tasks (Chang et al., 2024; Thirunavukarasu et al., 2023; Kasneci et al., 2023), but they may struggle with more complex reasoning problems (Zheng et al., 2023; Nezhurina et al., 2024). To further improve their reasoning ability, early works leverage prompting-based methods, such as Chain-of-Thought  (Wei et al., 2022), and Graph-of-Thought  (Besta et al., 2024), which encourage models to generate and organize intermediate reasoning steps during inference. Think-on-graph (ToG) (Sun et al., 2023) also follow this paradigm to prompt LLM reason on the graph. While effective, these methods are often sensitive to prompt design and computationally inefficient. Building on this, more recent approaches (Xu et al., 2025; Sui et al., 2025) focus on supervised fine-tuning (SFT) with curated reasoning traces (Deng et al., 2021; Lobo et al., 2024) and reinforcement learning (RL) with process or outcome rewards (Rafailov et al., 2023; Zhang et al., 2024). These methods allow models to internalize reasoning strategies rather than relying only on inference-time prompting. However, most of these advances focus on text-only tasks, whereas reasoning over graph-structured knowledge introduces additional challenges, requiring models to explore nodes, edges, and multi-hop paths. Although several works (Wang et al., 2023b; Luo et al., 2025; Yu et al., 2025) have attempted to adapt LLMs for graph reasoning tasks, most of them explicitly decompose complex queries into simpler sub-queries and rely on external retrievers. It remains underexplored how to enable LLMs to automatically explore graphs and generate answers in a unified manner.
In this paper, we follow previous studies (Luo et al., 2023; He et al., 2024; Sun et al., 2023) and leverage the Knowledge Graph Question Answering (KGQA) task to evaluate the effectiveness of graph reasoning, as it inherently involves multi-hop relational reasoning and explicit interaction with graph structures to obtain correct answers.
3. Preliminary
In this section, we first provide definitions of the key concepts and tasks used throughout the paper. We then analyze the retrievers adopted in current popular GraphRAG methods.
3.1. Definitions
Knowledge Graphs (KGs). Let $\mathcal{G}=\{\mathcal{V},\mathcal{R},\mathcal{E}\}$ denote a knowledge graph, where $\mathcal{V}$ is the set of entities (nodes), $\mathcal{R}$ is the set of relations, and $\mathcal{E}$ is the set of edges. Each edge $e\in\mathcal{E}$ represents a directed link between the head and tail entities $h,t\in\mathcal{V}$ under a relation $r\in R$, and can be denoted as a triple $(h,r,t)$. For a given node $v\in\mathcal{V}$, we use $\mathcal{N}_{v}$ to denote its neighboring entities, $\mathcal{R}_{v}$ to denote the associated relations and $\mathcal{E}_{v}$ to denote the connected edges of node $v$.
In this paper, we also augment the knowledge graph with inverse edges. Specifically, for each triple $(h,r,t)\in\mathcal{E}$, we add its inverse edge $(t,r_{\text{inverse}},h)$, following prior works (Kazemi and Poole, 2018; Dutt et al., 2022). This augmentation allows the model to explore the graph in both directions and has been shown to improve reasoning performance.
Knowledge Graph Question Answering (KGQA). KGQA is the task of answering natural language questions by reasoning over a knowledge graph. Given a question and a KG, $\mathcal{G}=\{\mathcal{V},\mathcal{R},\mathcal{E}\}$, the goal is to identify the correct entity or set of entities in $\mathcal{V}$ that answer the query. This typically requires multi-hop reasoning, where the model must traverse nodes and relations along valid paths (e.g., $(h,r_{1},v_{1}),(v_{1},r_{2},t)$) to reach the answer.
Due to the large size of knowledge graphs, which cannot fit into the limited context window of LLMs at once, most KGQA methods (i.e., GraphRAG) first retrieve a subgraph relevant to the query and then restrict the LLM’s reasoning to this subgraph. However, in this two-stage pipeline, the retriever and generator are usually optimized independently and executed sequentially. In the following subsections, we present experiments that illustrate the limitations of this design, focusing on the quality of the retrieved subgraphs, which can greatly affect the overall generation performance.
3.2. Experimental Settings
To conduct a comprehensive evaluation, we select one representative method from each retriever category as introduced in Section 2: G-Retriever (He et al., 2024) for heuristic-based methods, GNN-RAG (Mavromatis and Karypis, 2024) for GNN-based methods, and RoG (Luo et al., 2023) for LLM-based methods.
We select two widely used datasets: WebQSP (Yih et al., 2016), which primarily contains questions requiring one- or two-hop reasoning, and CWQ (Talmor and Berant, 2018), which consists of more complex questions that require multi-hop reasoning. For each question, there can be multiple correct answers. On average, the number of answers per test question is 10.20 for WebQSP and 1.89 for CWQ. For the heuristic-based method, i.e., G-Retriever, no training of the retriever is required. In contrast, both GNN-RAG and RoG require training. Specifically, for GNN-RAG, we train a GNN model as the retriever on each dataset, while for RoG we fine-tune the Llama-3.1-8B-Instruct model (Touvron et al., 2023) to serve as the retriever on each dataset. To evaluate retriever performance, we adopt three metrics: Hit, which measures whether at least one correct answer is included in the retrieved content; Recall, which measures the proportion of ground-truth answers retrieved for each question; and Precision, which measures the ratio of correctly retrieved entities among all retrieved entities.
(a) In-distribution performance comparison of three retrievers on WebQSP and CWQ datasets. Bars show Hit, Recall, and Precision, while dashed lines indicate ground-truth (GT) upper bounds.
(b) Cross-distribution generalization performance between CWQ and WebQSP. Solid bars denote transfer results, while transparent bars show in-distribution results.
Figure 1. Retrieval and generalization performance of different methods on WebQSP and CWQ datasets.
3.3. In-Distribution Retriever Performance
In this subsection, we aim to measure the performance of different retrievers when they are both trained and tested on the same dataset (i.e., WebQSP or CWQ). The performance of the three selected methods is shown in Figure 1(a), where we also include the ground-truth Hit and Recall as upper bounds, since the underlying graph does not contain all the answers for every question.
From the results in Figure 1(a), we observe that none of the retrievers achieve perfect performance, even in terms of Hit. For example, the Recall and Precision of G-Retriever on the CWQ dataset are only 38.65 and 5.14, respectively, indicating that many correct answers are missing while substantial noise is included in the retrieved subgraphs. This is because G-Retriever relies on embedding similarity to retrieve nodes, which is often insufficient for handling complex queries. This aligns with the well-known limitation of RAG methods in retrieving evidence for multi-hop queries (Shao et al., 2023; Tang and Yang, 2024), where iterative retrieval strategies are usually adopted by combining intermediate results from the generator. Among the three methods, GNN-RAG performs best, but it still shows around a 10% gap compared to the ground-truth Hit upper bound on the CWQ dataset. This demonstrates that the limitations of current retrievers will substantially constrain downstream generation performance.
3.4. Cross-Distribution Generalization
In Section 3.3, we mainly studied the retriever performance when training and testing on the same dataset. However, in real-world applications it is often impractical to train a separate retriever for every new graph. In this subsection, we therefore explore the generalization ability of pretrained retrievers. Instead of fine-tuning the retrievers for each dataset as in GNN-RAG (Mavromatis and Karypis, 2024) and RoG (Luo et al., 2023), we directly apply a retriever pretrained on one dataset to test on the other, in order to evaluate its cross-distribution generalization.
We first conduct experiments on the WebQSP and CWQ datasets, which are both built on Freebase KG (Bollacker et al., 2008) but differ substantially in question distribution. WebQSP consists of short, naturally collected questions that require 1-2-hop reasoning, whereas CWQ contains synthetically expanded questions with up to 4 hops and more complex linguistic structures such as conjunctions and comparatives. As shown in Figure 1(b), both RoG and GNN-RAG suffer a significant performance drop under the cross-distribution setting. For instance, the Hit of GNN-RAG decreases from 92.07 to 55.09, and its Recall drops from 85.43 to 31.68 on WebQSP when using the retriever trained on CWQ, performing even worse than the heuristic-based G-Retriever, which requires no retriever pretraining.
We further evaluate cross-graph generalization by applying retrievers trained on WebQSP to datasets with different knowledge graphs, including MetaQA (Zhang et al., 2018), which uses the WikiMovies KG, and PathQuestion (Zhou et al., 2018), which uses a Freebase subset with altered relation names. As shown in Table 1, RoG fails to retrieve any valid content, while GNN-RAG retrieves overly large subgraphs with very low precision.
Table 1. Cross-graph generalization performance.
Webqsp->MetaQA Webqsp->PathQuestion
Hit Recall Precision Hit Recall Precision
RoG 0 0 0 0 0 0
GNN-RAG 38.3 13.36 0.02 63.8 51.7 0.05
Overall, these results show that representative GraphRAG retrievers generalize poorly. We attribute this to two major factors:
  1. (1)
    Distribution gap in questions. Retrievers are typically trained to recognize the lexical and structural patterns of questions within a single dataset (e.g., WebQSP or CWQ). When the question style or reasoning composition changes, such as shifting from short, 1-2-hop natural questions to complex, synthetic multi-hop questions, the retriever fails to capture the correct reasoning paths and relevant entities. This suggests that retrievers overfit to dataset-specific surface forms rather than learning transferable reasoning patterns.
  2. (2)
    Distribution gap in graphs. Differences in the underlying knowledge graphs, such as relation vocabulary, and entity type distributions, further limit generalization. GNN-based retrievers depend on node and edge embeddings learned from a specific graph topology, while LLM-based retrievers (e.g., RoG) rely on relation paths that may not align across graphs. As a result, both classes of retrievers struggle to adapt when transferred to new graphs with different relation semantics or structural patterns.
3.5. Discussions
From Section 3.3, we observe that a substantial portion of answers cannot be retrieved by existing retrievers. Consequently, even strong generators fail to produce these correct answers since the necessary evidence is missing from the retrieved subgraphs. This highlights a key limitation of current two-stage GraphRAG paradigms—the retriever and generator operate independently and cannot interact with each other on the graph during reasoning.
The results in Section 3.4 further suggest that existing retrievers are over-specialized to their training datasets and knowledge graphs. They lack the capability to adjust their retrieval behavior based on the evolving environment or intermediate reasoning outcomes. This static retrieval behavior prevents effective adaptation to unseen question types and graph structures.
These insights motivates a new perspective on GraphRAG: rather than decoupling retrieval and generation or retrieving all potentially relevant content at once, the model should be able to interact with the whole graph step by step, selecting and exploring new nodes and relations as reasoning progresses for the question. Such a reasoning-driven retrieval process enables the model to dynamically incorporate feedback from intermediate steps, allowing it to generalize across diverse questions and graph structures by learning to reason and explore instead of memorizing fixed retrieval patterns.
4. Method
Figure 2. The framework of RoE. The model begins from the seed entity (green), incrementally expands to explored entities (yellow), while discovering answer entities (red). Stage 1 uses SFT to learn gold reasoning paths, and the trained model is then used as the initial model for Stage 2, where RL refines the exploration policy based on reward feedback.
Motivated by the limitations of existing GraphRAG methods identified in Section 3, we propose Reasoning by Exploration (RoE), a unified framework that integrates retrieval and generation within a single process, as illustrated in Figure 2. Unlike traditional two-stage GraphRAG pipelines, where a retriever extracts a static subgraph before reasoning begins, RoE enables a single LLM to actively explore the entire knowledge graph step by step.
4.1. Exploration Formulation
We formulate reasoning in RoE as an exploration process over a knowledge graph, where the LLM incrementally traverses the graph to construct reasoning paths that lead to correct answers. Formally, we formulate the exploration process as follows:
Given a question $q$ and a knowledge graph $\mathcal{G}=\{\mathcal{V},\mathcal{R},\mathcal{E}\}$, the model begins from one or more seed entities $\mathcal{V}_{0}$ that are linked to entities mentioned in the question. We expect the model to explore one-hop neighbors from the currently explored nodes at each step, progressively expanding the reasoning space. Specifically, at each step $d$, the model maintains a set of current reasoning paths:
$\mathcal{P}_{d}=\{p_{1}^{d},p_{2}^{d},\ldots,p_{n}^{d}\},$
where each path $p_{i}^{d}$ is defined as
$p_{i}^{d}=[v_{0},r_{1},v_{1},\ldots,r_{k},v_{k}],$
starting from a seed entity $v_{0}\in\mathcal{V}_{0}$ and expanding through a sequence of relations and entities. We define frontier nodes as
$\mathcal{F}_{d}=\{v_{k}\mid p_{i}^{d}\in\mathcal{P}_{d}\},$
representing the boundary entities that can be expanded at the next step. At step $d$, the model observes a state
$s_{d}=(q,\mathcal{P}_{d},\mathcal{N}_{\mathcal{F}_{d}}),$
where $\mathcal{N}_{\mathcal{F}_{d}}$ denotes neighbor nodes and edges of all frontier nodes. The model then predicts an action $a$ that consists of two lists:
  • Answers: $\mathcal{A}_{d}=[ans_{1}^{d},ans_{2}^{d},\ldots,ans_{m}^{d}]$, representing entities predicted as current answers at step $d$.
  • New Exploration Paths: $\mathcal{P}_{d+1}=\{p_{i}^{d}\oplus(v_{k},r_{k+1},v_{k+1})\mid p_{i}^{d}=[v_{0},r_{1},v_{1},\ldots,r_{k},v_{k}]\in\mathcal{P}_{d},(v_{k},r_{k+1},v_{k+1})\in\mathcal{E}\}$, where $\oplus$ denotes the extension of an existing path with a newly explored triplet.
The exploration continues until the model predicts that no new paths should be expanded (i.e., $\mathcal{P}_{d+1}=\emptyset$) or a predefined maximum number of steps $D_{\max}$ is reached.
The final outputs include the aggregated predicted answers
$\mathcal{A}=\bigcup_{d}\mathcal{A}_{d},$
which enables the model to find multiple answers through reasoning paths of different lengths.
Through this design, retrieval and generation become mutually informed and jointly optimized, enabling LLMs to construct reasoning paths that are both complete and contextually relevant. However, empowering LLMs to perform such structured graph exploration introduces several challenges. First, LLMs are pretrained primarily on unstructured text and therefore lack the inherent ability to navigate and reason over graphs (Dai et al., 2024). Second, naively training an LLM on exploration trajectories may cause it to memorize specific graph patterns rather than learning transferable exploration strategies. To address these challenges, we introduce a two-stage training strategy that equips the model with both graph exploration ability and strong generalization across different graphs.
4.2. Stage I: Supervised Fine-Tuning for Step-wise Exploration
In the first stage, we aim to equip the LLM with the basic ability to perform step-wise graph exploration. Specifically, we require the gold actions for each step, which include both the step-wise answers and the new exploration paths as defined in Section 4.1. Previous works (Zhang et al., 2022b; Sun et al., 2019) typically use the shortest path from the query entities to the answers to supervise their models. However, we argue that relying solely on the shortest path is overly restrictive. In real knowledge graphs, there may exist multiple valid reasoning paths connecting the same query entity to the correct answer; constraining training to only the shortest one discourages exploration and limits generalization. Moreover, the model does not have a global view of the entire graph at each step. When several neighboring entities share the same relation with the current frontier entity, selecting only a single target node can be ambiguous, especially when entity names are represented as IDs.
To address these issues, we construct the gold exploration paths by including all reasoning paths whose lengths are below a threshold, ensuring that the model learns to explore multiple plausible reasoning routes. Additionally, when one node is selected for expansion, other nodes connected to the frontier entity through the same relation are also treated as valid expansion candidates. For the answers at each step, we use the newly explored entities that correspond to the ground-truth answers as the gold answers. The detailed procedure is summarized in Algorithm 1 in Appendix A.
In the supervised fine-tuning (SFT) stage, each training instance is represented as a pair $(s_{d},a_{d}^{*})$ in $\mathcal{D}_{\text{SFT}}$, where $s_{d}$ denotes the state at step $d$ (including the question, current reasoning paths, and local neighborhood), and $a_{d}^{*}$ is the corresponding gold action constructed in Algorithm 1. At each step, the model takes $s_{d}$ as input and predicts action $a_{d}=(\mathcal{A}_{d},\mathcal{P}_{d})$, which includes the step-wise answers and the new exploration paths. The training objective is to maximize the likelihood of the gold action $a_{d}^{*}$ given its state, leading to the loss:
$\mathcal{L}_{\text{SFT}}=-\!\sum_{d}\log P(a_{d}^{*}\mid s_{d};\theta),$
where $\theta$ denotes the model parameters. This loss encourages the model to imitate the expert exploration behavior, deciding which nodes to expand and which entities to output as answers at each step, thereby acquiring the fundamental ability for structured graph exploration before reinforcement-learning refinement in Stage II.
4.3. Stage II: Reinforcement Learning for Exploration Optimization
Although the supervised fine-tuning (SFT) stage enables the model to imitate gold trajectories, it may cause the LLM to memorize fixed exploration patterns within a single dataset, limiting its ability to generalize to unseen graphs or question types. To address this, we introduce a reinforcement learning (RL) stage that allows the model to interact with the knowledge graph and refine its exploration strategy through feedback rather than imitation.
During this stage, we not only encourage the LLM to generate correct step-wise answers and exploration paths, but also aim to promote broader and more diverse exploration behaviors, enabling the model to discover additional valid reasoning trajectories and answers beyond those seen in the gold data. To enable such flexible and interpretable optimization, we adopt a rule-based reinforcement learning framework, which provides explicit feedback signals without relying on a learned value model. This choice is inspired by recent advances in reasoning LLMs, such as DeepSeek-R1 (Guo et al., 2025b), which have shown that rule-based rewards can effectively guide reasoning and exploration, while simplifying training compared to the original RLHF framework (Ouyang et al., 2022).
In RoE, we expect the model to generate outputs in the correct format, align its predictions with the gold answers and exploration paths, and simultaneously discover new valid answers and paths while penalizing hallucinated ones. Therefore, we define the following five rewards:
Format Reward. Different from previous methods that rely on a “thinking” process, we directly prompt the LLM to generate both the step-wise answers and exploration paths in a structured JSON format to improve efficiency. Specifically, at each step the model is required to output:
$\{\texttt{"answers": [\,], "exploration\_paths": [\,]}\}$.
A non-zero reward $\mathcal{R}_{\text{format}}$ is given only when the model successfully follows the required format. This encourages the LLM to produce well-structured outputs, facilitating reliable parsing and consistent downstream evaluation during training.
Answer Reward. We measure how many of the gold answers $\mathcal{A}_{d}^{*}$ in the current state $s_{d}$ are successfully predicted by the model. Specifically, we compute the recall between the predicted and gold answer sets at each step $d$:
$\mathcal{R}_{\text{ans}}=\frac{|\mathcal{A}_{d}^{p}\cap\mathcal{A}_{d}^{*}|}{|\mathcal{A}_{d}^{*}|},$
where $\mathcal{A}_{d}^{p}$ denotes the predicted answers. A higher reward is assigned when more ground-truth answers are correctly identified.
Answer Discovery Reward. To promote broader and more effective exploration, we further encourage the model to discover new valid answers that were not present in the previous steps. Let $\mathcal{A}$ denote the set of all gold answers for the question. We reward the model for predicting new correct answers $\mathcal{A}_{d}^{p}\cap(\mathcal{A}\setminus\mathcal{A}_{d}^{*})$ and penalize it for generating invalid ones $\mathcal{A}_{d}^{p}\setminus\mathcal{A}$:
$\mathcal{R}_{\text{ans-dis}}=|\mathcal{A}_{d}^{p}\cap(\mathcal{A}\setminus\mathcal{A}_{d}^{*})|-\beta\cdot|\mathcal{A}_{d}^{p}\setminus\mathcal{A}|,$
where $\beta$ is a penalty weight controlling the impact of hallucinated predictions. This reward encourages the model to find novel correct answers, by leveraging its internal knowledge and the reasoning paths explored so far, while maintaining answer precision.
Exploration Reward. Similar to the answer reward, this reward evaluates how well the model’s predicted reasoning paths match the gold paths in the current state. Let $\mathcal{P}_{d}^{p}$ and $\mathcal{P}_{d}^{*}$ denote the predicted and gold reasoning paths at step $d$, respectively. We compute the reward as the recall between the predicted and gold path sets:
$\mathcal{R}_{\text{explore}}=\frac{|\mathcal{P}_{d}^{p}\cap\mathcal{P}_{d}^{*}|}{|\mathcal{P}_{d}^{*}|}.$
A higher reward indicates that the model successfully explores the correct reasoning paths in the current step, guiding it to perform faithful and consistent graph traversal.
Exploration Discovery Reward. To further promote effective and diverse reasoning, we encourage the model to explore new valid paths while penalizing invalid ones. Let $\mathcal{P}_{d}^{p}$ denote the set of predicted paths at step $d$, and $\mathcal{P}_{d}^{*}$ denote the set of gold paths. We define new paths as those that do not appear in $\mathcal{P}_{d}^{*}$ but exist on the knowledge graph $\mathcal{G}$, and invalid paths as those whose triplets are not present in $\mathcal{G}$. The reward is computed as:
$\mathcal{R}_{\text{exp-dis}}=|\mathcal{P}_{d}^{p}\setminus\mathcal{P}_{d}^{*}|_{\text{valid}}-\beta\cdot|\mathcal{P}_{d}^{p}|_{\text{invalid}},$
where $|\mathcal{P}_{d}^{p}\setminus\mathcal{P}_{d}^{*}|_{\text{valid}}$ counts the number of new but valid paths on the graph and $|\mathcal{P}_{d}^{p}|_{\text{invalid}}$ counts the number of paths containing nonexistent triplets. This reward encourages the model to explore novel and meaningful reasoning directions while discouraging invalid graph traversals.
Final Reward. The overall reward for each step is defined as the sum of all the individual rewards introduced above:
$\mathcal{R}_{\text{total}}=\mathcal{R}_{\text{format}}+\mathcal{R}_{\text{ans}}+\mathcal{R}_{\text{ans-dis}}+\mathcal{R}_{\text{explore}}+\mathcal{R}_{\text{exp-dis}}.$
This combined reward jointly encourages the model to produce well-structured outputs, generate correct answers, and discover diverse and valid reasoning paths during reinforcement learning.
We optimize RoE using Group Relative Policy Optimization (GRPO) (Guo et al., 2025b), a variant of Proximal Policy Optimization (PPO) (Schulman et al., 2017). GRPO eliminates the need for an additional value function approximation by using the average reward of multiple sampled outputs as a relative baseline. For each input question $x$, GRPO samples a group of outputs $\{y_{1},y_{2},\ldots,y_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$ and optimizes the new policy $\pi_{\theta}$ by maximizing the following objective:
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{x\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\min\Big(r_{i,t}(\theta)\,\hat{A}_{i,t},$
$\quad\mathrm{clip}\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,\hat{A}_{i,t}\Big)-\gamma D_{\text{KL}}\!\big[\pi_{\theta}(\cdot|x)\,\|\,\pi_{\text{ref}}(\cdot|x)\big]\Bigg]$
where
$r_{i,t}(\theta)=\frac{\pi_{\theta}(y_{i,t}\mid x,y_{i,<t})}{\pi_{\theta_{\text{old}}}(y_{i,t}\mid x,y_{i,<t})},\qquad\hat{A}_{i,t}=\frac{R_{i}-\mathrm{mean}(\{R_{j}\}_{j=1}^{G})}{\mathrm{std}(\{R_{j}\}_{j=1}^{G})},$
and $\gamma$ is a hyperparameter balancing task-specific rewards and KL-divergence regularization. This formulation encourages the model to prefer outputs with higher relative rewards within each group, stabilizing training while improving sample efficiency. In our case, the rewards $R_{i}$ correspond to the total rule-based rewards $\mathcal{R}_{\text{total}}$.
In summary, RoE is first supervised through step-wise fine-tuning (Stage I) to learn the fundamental exploration behavior from gold trajectories, and is then refined via reinforcement learning (Stage II) to improve its reasoning and generalization ability. Through this two-stage training framework, RoE learns to jointly reason and explore over graphs, constructing accurate and diverse reasoning paths while producing faithful multi-hop answers.
5. Experiments
In this section, we conduct comprehensive experiments to validate the effectiveness of the proposed RoE. Specifically, we aim to answer the following research questions: (1) RQ1: How does RoE perform on benchmark datasets compared with baseline methods? (2) RQ2: Can RoE generalize to unseen datasets better than existing GraphRAG approaches? (3) RQ3: How do different design choices and factors influence the performance of RoE?
Table 2. Performance comparison on two KGQA benchmarks. The best and second-best are in bold and underline.
indicates the model uses Llama3.1-8B-Instruct as backbone.
Category Method WebQSP CWQ Avg. Rank
Hits@1 F1 Hits@1 F1
LLMs-only Flan-T5-xl 31.00 - 14.70 - 12.00
Alpaca-7B 51.80 - 27.40 - 11.00
Llama-2-7B-Chat 64.40 - 34.60 - 8.50
Llama-3.1-8B-Instruct 58.96 31.66 31.49 20.66 10.25
GPT-4.1 64.99 44.04 44.09 36.56 8.25
GNNs-only GraftNet - 62.40 - 32.70 8.00
SR+NSM - 64.10 - 47.10 5.50
UniKGQA - 70.20 - 48.00 3.50
LLMs+GNNs G-Retriver 74.07 54.51 51.51 45.18 6.75
GNNRAG 85.25 72.01 63.61 54.92 2.00
LLMs+KGs ToG + Llama2-70B-Chat 63.70 - 53.60 - 7.50
RoG 78.62 64.39 53.83 45.97 5.00
KD-CoT 68.60 52.50 55.70 - 6.33
SubGraphRAG 86.61 70.57 56.98 47.16 3.00
RoE 89.13 74.76 66.49 53.21 1.25
5.1. Experimental Settings
Datasets. To evaluate the effectiveness of RoE, we use two widely adopted datasets: WebQSP (Yih et al., 2016) and CWQ (Talmor and Berant, 2018), both constructed upon the Freebase knowledge graph (Bollacker et al., 2008). To further assess the generalization ability of GraphRAG methods across different knowledge graphs, we additionally use MetaQA (Zhang et al., 2018), which is built on the WikiMovies KG (Miller et al., 2016), and PathQuestion (Zhou et al., 2018), which uses a Freebase subset with modified relation names. For cross-graph generalization evaluation, we randomly sample 1,000 multi-hop questions from each of these two datasets.
Evaluation Metrics. We follow previous works (He et al., 2021; Guo et al., 2025a; Li et al., 2024) and adopt two evaluation metrics: Hit and F1. Specifically, the Hit measures the proportion of questions for which at least one correct predicted answer, while F1 provides a balanced evaluation of both precision and recall.
Baselines. We compare RoE with a diverse set of baselines, which can be grouped into four categories: (1) LLM-Only: methods that rely solely on the internal knowledge of LLMs to answer questions without using external knowledge graphs. We include Flan-T5-XL (Chung et al., 2024), Alpaca-7B (Taori et al., 2023), Llama-2-7B-Chat (Touvron et al., 2023), Llama-3.1-8B-Instruct (Touvron et al., 2023), and GPT-4.1 (Achiam et al., 2023). (2) GNN-Only: models that use GNNs for both retrieval and reasoning. We include GraftNet (Sun et al., 2018), SR+NSM (Zhang et al., 2022b), and UniKGQA (Jiang et al., 2022). (3) LLMs + GNNs: hybrid approaches that combine LLMs and GNNs for retrieval and generation. For instance, G-Retriever (He et al., 2024) leverages GNNs to enhance LLM reasoning, while GNN-RAG (Mavromatis and Karypis, 2024) employs GNNs for retrieval and LLMs for generation. (4) LLMs + KG: methods that employ LLMs for both retrieval and generation over knowledge graphs. We compare with ToG (Sun et al., 2023), RoG (Luo et al., 2023), KD-CoT (Wang et al., 2023a), and SubGraphRAG (Li et al., 2024).
RoE settings. For RoE, we use the Llama-3.1-8B-Instruct model (Touvron et al., 2023) as the backbone. We divide the training data of both WebQSP and CWQ into two parts: 60% is used for the SFT (stage I) and the remaining 40% for the RL (stage II). To reduce memory consumption during training, we apply the LoRA (Hu et al., 2022) in both stages.
To ensure a fair comparison, we train each retriever and generator on a single dataset, rather than jointly training across multiple datasets as done in RoG and GNN-RAG. Moreover, to eliminate the influence of backbone model, we use the same Llama-3.1-8B-Instruct model (Touvron et al., 2023) for most LLM-based baselines, including RoG, KD-CoT, SubGraphRAG, G-Retriever, and GNN-RAG.
5.2. Overall Performance Comparison
The overall performance on the WebQSP and CWQ datasets is shown in Table 2. We make the following observations:
  • The proposed RoE achieves the best performance across most metrics on both datasets. Specifically, RoE attains relative improvements of 2.9% and 3.8% on Hit and F1, respectively, compared to the second-best models on the WebQSP dataset. On the CWQ dataset, RoE also achieves a 3.1% relative improvement in Hit, demonstrating its strong ability to reason on graphs.
  • The LLM-only models, which lack access to external graph information, generally perform worse than methods that incorporate graphs. However, stronger models, such as GPT-4.1, achieve better results, suggesting that large LLMs already possess internal knowledge relevant to the benchmarks. This observation justifies our choice that using the same LLM backbone for baselines to ensure a fair comparison.
  • The GNN-only models generally perform worse than methods that incorporate LLMs and graph information, indicating that the internal knowledge and reasoning capability of LLMs are essential for solving multi-hop KGQA tasks.
  • Compared with LLMs+GNNs methods, RoE achieves better performance in most cases, indicating additional GNN modules may not be necessary, suggesting LLMs themselves can acquire effective graph reasoning capabilities through fine-tuning.
  • For the LLMs+KGs methods, such as ToG, KD-CoT, and SubGraphRAG, which prompt pretrained LLMs to iteratively retrieve and generate sub-questions or directly answer based on a retrieved subgraph, the performance is still lower than that of the proposed RoE. This suggests that pretrained LLMs are not well-suited for graph reasoning without additional fine-tuning.
5.3. Generalization Performance Comparison
(a) CWQ → WebQSP
(b) WebQSP → MetaQA
(c) WebQSP → PathQuestion
Figure 3. Generalization performance (Hit/F1) of different methods across dataset transfers.
In this section, we evaluate the generalization capability of RoE. Unlike previous works (Luo et al., 2023; Mavromatis and Karypis, 2024), which fine-tune their models on the target graphs, we directly apply the pretrained model from one dataset to unseen datasets without any additional training. Specifically, we select G-Retriever, RoG, and GNN-RAG as baselines, and evaluate the transfer performance under three settings: CWQ $\rightarrow$ WebQSP, WebQSP $\rightarrow$ PathQuestion, and WebQSP $\rightarrow$ MetaQA.
The generalization results are presented in Figure 3. As shown, the proposed RoE significantly outperforms all baselines across different transfer settings. For example, under the WebQSP $\rightarrow$ MetaQA transfer, RoE achieves a Hit of 93.90, while the second-best method only reaches 52.30, demonstrating the strong generalization capability of RoE across distributional shifts in both questions and knowledge graphs. Specifically, although G-Retriever employs a heuristic-based retriever without explicit retriever training, it still requires training a GNN model to encode the graph structure. As a result, G-Retriever performs well on the CWQ $\rightarrow$ WebQSP transfer setting, where both datasets share the same underlying knowledge graph, but performs poorly in other settings with different graph structures. Similarly, the RoG model relies on predicting reasoning relation paths, which are difficult to transfer across graphs with different relation vocabularies. GNN-RAG relies on trained GNNs for retrieval, which also makes it difficult to generalize to different question distributions and knowledge graphs. As a result, it tends to retrieve overly large subgraphs under transfer settings, as discussed in Section 3, thereby introducing substantial noise that negatively affects the generator’s reasoning performance.
5.4. Ablation Studies
In this subsection, we conduct ablation studies to verify the effectiveness of the key components in RoE. Specifically, we analyze the impact of the reinforcement learning (RL) stage and the proposed discovery rewards for both answers and exploration paths. We conduct experiments on two variants of RoE: (1) removing the reinforcement learning (RL) stage and using only supervised fine-tuning (SFT), denoted as RoE w/o RL; and (2) removing the discovery rewards for both Answer Discovery Reward and Exploration Discovery Reward during the RL stage, denoted as RoE w/o Discovery. We train these variants on WebQSP dataset and test their performance on WebQSP, MetaQA and PathQuestion dataset.
Figure 4. Performance of RoE and its variants pretrained on the WebQSP dataset. Marker shapes denote different model variants, while filled and hollow markers represent Hit and metrics, respectively.
The results are presented in Figure 4, from which we make the following observations: (1) The RL stage is essential for RoE, especially under transfer settings. RoE w/o RL shows significantly lower performance on the MetaQA and PathQuestion datasets, demonstrating that models trained only with SFT tend to memorize dataset-specific patterns and struggle to generalize to unseen graphs without reinforcement learning. (2) The discovery rewards, which guide the model to find more valid answers and paths while penalizing hallucinations, are also crucial for performance. Without these rewards, the model tends to generate more hallucinated answers, resulting in a notable drop in F1 under transfer settings.
6. Conclusion
In this paper, we revisited the limitations of existing two-stage GraphRAG frameworks, where the retriever and generator are optimized independently and lack interaction during reasoning. We proposed Reasoning by Exploration (RoE), a unified framework that integrates retrieval and generation into a step-wise exploration process. By allowing the model to interact with the graph dynamically and incorporate feedback from intermediate steps, RoE learns to reason and explore rather than rely on static retrieval patterns. Extensive experiments on multiple KGQA benchmarks demonstrate that RoE consistently outperforms strong baselines and generalizes effectively across different question distributions and graphs.
Appendix A Algorithm of SFT Dataset Construction
In this section, we detail the algorithm to construct the SFT datasets to train RoE. The algorithm is summarized in Algorithm 1, where lines 1–14 construct all gold-consistent reasoning paths up to length $L_{\max}$; line 20 derives the step-wise gold answers; and lines 21–22 generate the step-wise new exploration paths.
Appendix B Datasets
In the experiments, we adopt four different datasets, i.e. WebQSP (Yih et al., 2016) and CWQ (Talmor and Berant, 2018), MetaQA (Zhang et al., 2018) and PathQuestion (Zhou et al., 2018). For the WebQSP and CWQ dataset, we leverage the processed dataset in Luo et al. (2023). For the PathQuestion and MetaQA dataset, we randomly sample 1,000 samples of two hop questions. The statistics of these datasets are shown in Table 3.
Table 3. Statistics of datasets.
Datasets #Train #Test Max #hop
WebQSP 2,826 1,628 2
CWQ 27,639 3,531 4
PathQuestion 0 1,000 2
MetaQA 0 1,000 2
Appendix C Experimental Settings
In this section, we present the detailed experimental settings for RoE. For RoE and most of baseline models involving LLMs, we use the Llama-3.1-8B-Instruct model (Touvron et al., 2023) as the backbone. We first leverage Algorithm 1 to construct the gold trajectories for supervised fine-tuning. The training data of both WebQSP and CWQ are divided into two parts: 60% for the Supervised Fine-Tuning (SFT) stage and the remaining 40% for the Reinforcement Learning (RL) stage. To reduce memory consumption, we adopt LoRA (Hu et al., 2022) in both stages with a rank of 32.
During SFT, the loss is calculated only on the action prediction component, and the model is optimized using the AdamW optimizer with a learning rate of $1\times 10^{-4}$. For the RL stage, we apply a smaller learning rate of $5\times 10^{-6}$ and use the Group Relative Policy Optimization (GRPO) algorithm to optimize the policy with rule-based rewards. For each sample, we generate 4 responses. Each training step uses a batch size of 4 with gradient accumulation to fit within GPU memory. Training is performed on $\timesH2$00 GPUs.
During inference, we follow a depth-first exploration strategy. Because each node in the knowledge graph may have numerous neighbors that cannot fit into the LLM context window at once, we split the neighbors into multiple batches and iteratively feed them to the model. We use a depth-first search algorithm to explore the graph, and we set the maximum exploration depth to 5 hops to prevent overly long reasoning chains and context overflow. For reproducibility, all random seeds are fixed, and the same prompt templates for SFT and RL.
Algorithm 1 SFT Dataset Construction
1:Question $q$; KG $\mathcal{G}=\{\mathcal{V},\mathcal{R},\mathcal{E}\}$; seed entities $\mathcal{V}_{0}$; gold answers $\mathcal{A}^{*}$; max path length $L_{\max}$
2:Dataset $\mathcal{D}_{\text{SFT}}=\{(s_{d},\ a_{d}^{*})\}_{d=0}^{L_{\max}-1}$ with $a_{d}^{*}=(\mathcal{A}_{d}^{*},\mathcal{P}_{d}^{*})$
3:Phase I: Mine all gold-consistent paths (length $\leq L_{\max}$)
[…]
5:while $\mathsf{Q}$ not empty do
6:  $p\leftarrow$ pop from $\mathsf{Q}$; $v\leftarrow\text{frontier}(p)$
7:  if $|p|-1>L_{\max}$ then continue
8:  end if
9:  if $v\in\mathcal{A}^{*}$ then $\Pi\leftarrow\Pi\cup\{p\}$
10:  end if
11:  for all $(v,r,u)\in\mathcal{E}$ do
12:   if $u\notin p$ then $\triangleright$ simple-path constraint (avoid cycles)
13:     $\mathsf{Q}\leftarrow\mathsf{Q}\cup\{\,p\oplus(v,r,u)\,\}$
14:   end if
15:  end for
16:end while
17:Phase II: Build step-wise actions from mined paths
18:Define depth-indexed prefix pools for $d=0,\dots,L_{\max}$:
$\mathsf{Pref}(d)\ =\ \{\ \text{all prefixes of any }p\in\Pi\text{ with }|{\rm prefix}|-1=d\ \}$
[…]
20:for $d=0$ to $L_{\max}-1$ do
21:  $\mathcal{P}_{d}\leftarrow\mathsf{Pref}(d)$; $\mathcal{F}_{d}\leftarrow\{\text{frontier}(p)\mid p\in\mathcal{P}_{d}\}$
[…]
23:  Gold step-answers:
$\mathcal{A}_{d}^{*}\ =\ \{\,u\in\mathcal{A}^{*}\ \mid\ \exists(v,r,u)\in\mathcal{N}_{\mathcal{F}_{d}}\,\}$
24:  Gold one-hop extensions:
$\mathcal{E}^{\Pi}_{d}\ =\ \{\,(p,(v,r,u))\in\mathsf{Pref}(d+1)\ \ \mid\ p\in\mathcal{P}_{d},\ \text{frontier}(p)=v\}$
25:  Same-relation sibling expansion:
$\mathcal{P}_{d}^{*}\ =\ \big\{\,p\oplus(v,r,u^{\prime})\ \big|\ (p,(v,r,u))\in\mathcal{E}^{\Pi}_{d},\ (v,r,u^{\prime})\in\mathcal{E}\big\}$
26:  Record step: $a_{d}^{*}\leftarrow(\mathcal{A}_{d}^{*},\mathcal{P}_{d}^{*})$; $\mathcal{D}_{\text{SFT}}\leftarrow\mathcal{D}_{\text{SFT}}\cup\{(s_{d},\ a_{d}^{*})\}$
27:end for
28:return $\mathcal{D}_{\text{SFT}}$

### v2

@@
 Experiments on benchmark datasets demonstrate that RoE achieves substantial overall improvements over baselines, while also generalizing effectively to unseen graphs.
+The code can be found at https://github.com/haoyuhan1/RoE.
 Large Language Models, Graph Reasoning, Retrieval-Augmented Generation, Multi-hop Question Answering
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NY
+journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792684isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Language modelsccs: Information systems Question answeringccs: Computing methodologies Reinforcement learning
 1. Introduction
 Large language models (LLMs) (Chang et al., 2024; Thirunavukarasu et al., 2023; Kasneci et al., 2023) have demonstrated impressive abilities across a wide range of natural language processing tasks.
-To further enhance their reliability and factuality, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Gao et al., 2023; Jiang et al., 2023) has become a widely adopted paradigm.
+To further enhance their reliability and factuality, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Gao et al., 2023; Jiang et al., 2023; Wang et al., 2024) has become a widely adopted paradigm.
 By retrieving external knowledge before generating responses, RAG has achieved strong performance in diverse domains, such as healthcare (Xu et al., 2024), law (Wiratunga et al., 2024), finance (Zhang et al., 2023), and education (Chu et al., 2025).
@@
 Graphs provide a natural representation for structured knowledge, capturing rich relational dependencies among entities.
-They are widely used in domains such as knowledge bases, social networks, and scientific applications including biology and chemistry (Wu et al., 2020; Ma and Tang, 2021; Wu et al., 2023a).
-Recently, Graph Retrieval-Augmented Generation (GraphRAG) has emerged as a promising way to integrate such graph-structured knowledge into LLMs (Han et al., 2024; Peng et al., 2024).
-There are mainly two steps in GraphRAG: first, a retriever selects query-relevant content from the entire graph, where the retrieved units may be nodes, paths, or subgraphs; then, a generator, usually an LLM, produces answers conditioned on the retrieved subgraphs.
-This framework has been successfully applied to tasks such as knowledge graph question answering (KGQA) (Yasunaga et al., 2021, 2022; Zhang et al., 2022a), where answering a query often requires multi-step reasoning over the graphs.
+They are widely used in domains such as knowledge bases, social networks, and scientific applications including biology and chemistry (Wu et al., 2020; Ma and Tang, 2021; Wang et al., 2020).
+Recently, Graph Retrieval-Augmented Generation (GraphRAG) has emerged as a promising way to integrate such graph-structured knowledge into LLMs (Han et al., 2024; Peng et al., 2024; Lei et al., 2025b).
+There are mainly two steps in GraphRAG: first, a retriever selects query-relevant content from the entire graph, where the retrieved units may be nodes, paths, or subgraphs; and then, a generator, usually an LLM, produces answers conditioned on the retrieved subgraphs.
+This framework has been successfully applied to tasks such as knowledge graph question answering (KGQA) (Yasunaga et al., 2021, 2022; Zhang et al., 2022b), where often requires multi-step reasoning over the graphs.
 While efficient, this two-phase pipeline may constrain reasoning quality due to the disjoint optimization of retriever and generator.
@@
 This query may have multiple answers and can be reached through multiple reasoning paths, such as “Churchill $\rightarrow$ sibling $\rightarrow$ child” or “Churchill $\rightarrow$ parent $\rightarrow$ child $\rightarrow$ child”.
-Heuristic-based retrievers (Yasunaga et al., 2022; Feng et al., 2020) (e.g., retrieving fixed $k$-hop subgraphs) or GNN-based retrievers (Mavromatis and Karypis, 2024; Liu et al., 2024) may include a large amount of irrelevant information about Churchill, overwhelming the generator with noise (Guo et al., 2025a).
-In contrast, LLM-based retrievers (Luo et al., 2023; Wu et al., 2023b) may generate only a limited set of reasoning paths, leading to missing valid entities and incomplete answers, as shown in Section 3.
-Second, retrievers or generators trained on specific graphs or datasets usually generalize poorly to unseen graphs as shown in Section 3.
-For instance, a path-based retriever tuned on a movie knowledge base may learn to generate reasoning paths primarily around relations such as “actor $\rightarrow$ film $\rightarrow$ director”, but it struggles to generalize when applied to graphs with very different relational structures, such as biomedical graphs involving “gene $\rightarrow$ protein $\rightarrow$ disease” paths.
+Heuristic-based retrievers (Yasunaga et al., 2022; Feng et al., 2020) (e.g., retrieving fixed $k$-hop subgraphs) or GNN-based retrievers (Mavromatis and Karypis, 2024; Liu et al., 2024) may include a large amount of irrelevant information about Churchill, overwhelming the generator with noise (Guo et al., 2025b).
+In contrast, LLM-based retrievers (Luo et al., 2023; Wu et al., 2023) may generate only a limited set of reasoning paths, leading to missing valid entities and incomplete answers, as shown in Section 3.
+Second, retrievers or generators trained on specific graphs or datasets usually generalize poorly to unseen graphs and question types as shown in Section 3.
+For instance, a path-based retriever tuned on a movie knowledge base may learn to generate reasoning paths primarily around relations such as “actor $\rightarrow$ film $\rightarrow$ director”, but it may struggle to generalize when applied to graphs in other domains with very different relational structures, such as biomedical graphs involving “gene $\rightarrow$ protein $\rightarrow$ disease” paths.
 This dependence leads to sharp performance drops when deployed on unseen graphs, limiting the robustness and real-world applicability of current GraphRAGs.
@@
 In the first stage, we construct gold exploration trajectories and apply supervised fine-tuning (SFT) to teach the LLM to expand nodes and edges step by step.
-In the second stage, we employ reinforcement learning (RL) with different reward signals to further refine the exploration policy, encouraging the model to efficiently discover valid reasoning paths while improving its ability to generalize across diverse and unseen graphs.
-This combination enables RoE to leverage both explicit supervision and adaptive learning signals, resulting in more faithful reasoning and stronger generalization to unseen graphs.
-Our key contributions are summarized as follows:
-  • We introduce Reasoning by Exploration (RoE), a unified framework that integrates retrieval and generation into a single process by framing reasoning as graph exploration, allowing LLMs to construct reasoning paths and answers simultaneously.
-  • We develop a two-stage training strategy that equips LLMs with exploration ability: supervised fine-tuning (SFT) on gold reasoning trajectories to learn step-wise expansion, followed by reinforcement learning (RL) to refine exploration policies and enhance generalization to unseen graphs.
-  • We conduct extensive experiments on multi-hop reasoning benchmarks, demonstrating that RoE significantly outperforms state-of-the-art baselines while exhibiting strong generalization across diverse unseen graphs.
+In the second stage, we employ reinforcement learning (RL) to further refine the exploration policy, encouraging the model to efficiently discover valid reasoning paths while improving its ability to generalize across diverse and unseen graphs.
+Together, this training paradigm enables RoE to combine explicit supervision with adaptive learning signals, leading to more faithful reasoning and stronger generalization.
+We conduct extensive experiments on multiple multi-hop reasoning benchmarks, showing that RoE consistently outperforms strong state-of-the-art baselines and exhibits robust generalization to unseen graphs and question types across different backbone sizes.
 2. Related Works
 2.1. GraphRAG
-GraphRAG (Han et al., 2024; Peng et al., 2024) aims to retrieve relevant information from external graphs to improve the performance and reduce the hallucination of LLMs.
+GraphRAG (Han et al., 2024; Peng et al., 2024; Lei et al., 2025a; Han et al., 2025a) aims to retrieve relevant information from external graphs to improve the performance and reduce the hallucination of LLMs.
 Compared to traditional RAG systems (Lewis et al., 2020; Gao et al., 2023) that operate on unstructured text, GraphRAG leverages the relational structure of graphs, where nodes represent entities and edges encode relations between nodes.
@@
 These seeds are then expanded by retrieving their $k$-hop neighbors (Yasunaga et al., 2021; Choudhary and Reddy, 2023), forming a candidate subgraph for reasoning.
-A representative example is G-Retriever (He et al., 2024), which first leverages embedding models to retrieve the most similar nodes and edges to the query, and then applies the Prize-Collecting Steiner Tree (PCST) algorithm to extract a compact subgraph centered around the selected seeds.
+A representative example is G-Retriever (He et al., 2024), which first leverages embedding models to retrieve the most similar nodes and edges to the query, and then applies the Prize-Collecting Steiner Tree (PCST) algorithm to extract a compact subgraph.
 GNN-based Retriever.
@@
 For example, GNN-RAG (Mavromatis and Karypis, 2024) first trains a GNN model and then retrieves the nodes with the highest probability scores as candidate answers, while also identifying the shortest paths connecting the question entities with these candidates to construct supporting evidence.
-Compared to heuristic-based methods, GNN-based retrievers can capture richer relational dependencies and are more adaptive to complex queries.
 LLM-based Retriever.
@@
 2.2. Reasoning of Large Language Models
-Large Language Models (LLMs) have demonstrated strong performance on natural language tasks (Chang et al., 2024; Thirunavukarasu et al., 2023; Kasneci et al., 2023), but they may struggle with more complex reasoning problems (Zheng et al., 2023; Nezhurina et al., 2024).
-To further improve their reasoning ability, early works leverage prompting-based methods, such as Chain-of-Thought  (Wei et al., 2022), and Graph-of-Thought  (Besta et al., 2024), which encourage models to generate and organize intermediate reasoning steps during inference.
+Large Language Models (LLMs) have demonstrated strong performance on natural language tasks (Chang et al., 2024; Thirunavukarasu et al., 2023; Kasneci et al., 2023; Han et al., 2025b), but they may struggle with more complex reasoning problems (Zheng et al., 2023; Nezhurina et al., 2024).
+To further improve their reasoning ability, early works leverage prompting-based methods, such as Chain-of-Thought  (Wei et al., 2022), Tree-of-Thought (Yao et al., 2023) and Graph-of-Thought  (Besta et al., 2024), which encourage models to generate and organize intermediate reasoning steps during inference.
 Think-on-graph (ToG) (Sun et al., 2023) also follow this paradigm to prompt LLM reason on the graph.
@@
 However, most of these advances focus on text-only tasks, whereas reasoning over graph-structured knowledge introduces additional challenges, requiring models to explore nodes, edges, and multi-hop paths.
-Although several works (Wang et al., 2023b; Luo et al., 2025; Yu et al., 2025) have attempted to adapt LLMs for graph reasoning tasks, most of them explicitly decompose complex queries into simpler sub-queries and rely on external retrievers.
+Although several works (Wang et al., 2023a; Luo et al., 2025; Yu et al., 2025) have attempted to adapt LLMs for graph reasoning tasks, most of them explicitly decompose complex queries into simpler sub-queries and rely on external retrievers.
 It remains underexplored how to enable LLMs to automatically explore graphs and generate answers in a unified manner.
@@
 3. Preliminary
-In this section, we first provide definitions of the key concepts and tasks used throughout the paper. We then analyze the retrievers adopted in current popular GraphRAG methods.
+In this section, we first provide definitions, and then analyze the retrievers adopted in current popular GraphRAG methods.
 3.1. Definitions
 Knowledge Graphs (KGs).
-Let $\mathcal{G}=\{\mathcal{V},\mathcal{R},\mathcal{E}\}$ denote a knowledge graph, where $\mathcal{V}$ is the set of entities (nodes), $\mathcal{R}$ is the set of relations, and $\mathcal{E}$ is the set of edges.
-Each edge $e\in\mathcal{E}$ represents a directed link between the head and tail entities $h,t\in\mathcal{V}$ under a relation $r\in R$, and can be denoted as a triple $(h,r,t)$.
-For a given node $v\in\mathcal{V}$, we use $\mathcal{N}_{v}$ to denote its neighboring entities, $\mathcal{R}_{v}$ to denote the associated relations and $\mathcal{E}_{v}$ to denote the connected edges of node $v$.
+Let $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{E})$ be a knowledge graph, where $\mathcal{V}$ denotes entities, $\mathcal{R}$ relation types, and $\mathcal{E}$ directed edges.
+Each edge is represented as a triple $(h,r,t)$ with $h,t\in\mathcal{V}$ and $r\in\mathcal{R}$.
+For any node $v\in\mathcal{V}$, we denote its neighboring entities, incident relations, and connected edges as $\mathcal{N}_{v}$, $\mathcal{R}_{v}$, and $\mathcal{E}_{v}$, respectively.
 In this paper, we also augment the knowledge graph with inverse edges.
 Specifically, for each triple $(h,r,t)\in\mathcal{E}$, we add its inverse edge $(t,r_{\text{inverse}},h)$, following prior works (Kazemi and Poole, 2018; Dutt et al., 2022).
-This augmentation allows the model to explore the graph in both directions and has been shown to improve reasoning performance.
 Knowledge Graph Question Answering (KGQA).
@@
 The performance of the three selected methods is shown in Figure 1(a), where we also include the ground-truth Hit and Recall as upper bounds, since the underlying graph does not contain all the answers for every question.
-From the results in Figure 1(a), we observe that none of the retrievers achieve perfect performance, even in terms of Hit.
+From the results, we observe that none of the retrievers achieve perfect performance.
 For example, the Recall and Precision of G-Retriever on the CWQ dataset are only 38.65 and 5.14, respectively, indicating that many correct answers are missing while substantial noise is included in the retrieved subgraphs.
-This is because G-Retriever relies on embedding similarity to retrieve nodes, which is often insufficient for handling complex queries.
-This aligns with the well-known limitation of RAG methods in retrieving evidence for multi-hop queries (Shao et al., 2023; Tang and Yang, 2024), where iterative retrieval strategies are usually adopted by combining intermediate results from the generator.
 Among the three methods, GNN-RAG performs best, but it still shows around a 10% gap compared to the ground-truth Hit upper bound on the CWQ dataset.
@@
 $\mathcal{P}_{d}=\{p_{1}^{d},p_{2}^{d},\ldots,p_{n}^{d}\},$
-where each path $p_{i}^{d}$ is defined as
-$p_{i}^{d}=[v_{0},r_{1},v_{1},\ldots,r_{k},v_{k}],$
-starting from a seed entity $v_{0}\in\mathcal{V}_{0}$ and expanding through a sequence of relations and entities. We define frontier nodes as
+where each path $p_{i}^{d}$ is defined as $p_{i}^{d}=[v_{0},r_{1},v_{1},\ldots,r_{k},v_{k}],$ starting from a seed entity $v_{0}\in\mathcal{V}_{0}$ and expanding through a sequence of relations and entities.
+We define frontier nodes as
 $\mathcal{F}_{d}=\{v_{k}\mid p_{i}^{d}\in\mathcal{P}_{d}\},$
@@
 The exploration continues until the model predicts that no new paths should be expanded (i.e., $\mathcal{P}_{d+1}=\emptyset$) or a predefined maximum number of steps $D_{\max}$ is reached.
-The final outputs include the aggregated predicted answers
-$\mathcal{A}=\bigcup_{d}\mathcal{A}_{d},$
-which enables the model to find multiple answers through reasoning paths of different lengths.
+The final outputs include all predicted answers $\mathcal{A}=\bigcup_{d}\mathcal{A}_{d}$, which enables the model to find multiple answers through reasoning paths of different lengths.
 Through this design, retrieval and generation become mutually informed and jointly optimized, enabling LLMs to construct reasoning paths that are both complete and contextually relevant.
@@
 Specifically, we require the gold actions for each step, which include both the step-wise answers and the new exploration paths as defined in Section 4.1.
-Previous works (Zhang et al., 2022b; Sun et al., 2019) typically use the shortest path from the query entities to the answers to supervise their models.
+Previous works (Zhang et al., 2022a; Sun et al., 2019) typically use the shortest path from the query entities to the answers to supervise their models.
 However, we argue that relying solely on the shortest path is overly restrictive.
@@
 At each step, the model takes $s_{d}$ as input and predicts action $a_{d}=(\mathcal{A}_{d},\mathcal{P}_{d})$, which includes the step-wise answers and the new exploration paths.
-The training objective is to maximize the likelihood of the gold action $a_{d}^{*}$ given its state, leading to the loss:
+The training objective is to maximize the likelihood of the gold action $a_{d}^{*}$ given its state, leading to loss:
 $\mathcal{L}_{\text{SFT}}=-\!\sum_{d}\log P(a_{d}^{*}\mid s_{d};\theta),$
 where $\theta$ denotes the model parameters.
-This loss encourages the model to imitate the expert exploration behavior, deciding which nodes to expand and which entities to output as answers at each step, thereby acquiring the fundamental ability for structured graph exploration before reinforcement-learning refinement in Stage II.
+This loss encourages the model to imitate the expert exploration behavior, thereby acquiring the fundamental ability for structured graph exploration before reinforcement-learning refinement in Stage II.
 4.3. Stage II: Reinforcement Learning for Exploration Optimization
-Although the supervised fine-tuning (SFT) stage enables the model to imitate gold trajectories, it may cause the LLM to memorize fixed exploration patterns within a single dataset, limiting its ability to generalize to unseen graphs or question types.
+Although the SFT stage enables the model to imitate gold trajectories, it may cause the LLM to memorize fixed exploration patterns within a single dataset, limiting its ability to generalize to unseen graphs or question types.
 To address this, we introduce a reinforcement learning (RL) stage that allows the model to interact with the knowledge graph and refine its exploration strategy through feedback rather than imitation.
@@
 To enable such flexible and interpretable optimization, we adopt a rule-based reinforcement learning framework, which provides explicit feedback signals without relying on a learned value model.
-This choice is inspired by recent advances in reasoning LLMs, such as DeepSeek-R1 (Guo et al., 2025b), which have shown that rule-based rewards can effectively guide reasoning and exploration, while simplifying training compared to the original RLHF framework (Ouyang et al., 2022).
+This choice is inspired by recent advances in reasoning LLMs, such as DeepSeek-R1 (Guo et al., 2025a), which have shown that rule-based rewards can effectively guide reasoning and exploration, while simplifying training compared to the original RLHF framework (Ouyang et al., 2022).
 In RoE, we expect the model to generate outputs in the correct format, align its predictions with the gold answers and exploration paths, and simultaneously discover new valid answers and paths while penalizing hallucinated ones.
@@
 A non-zero reward $\mathcal{R}_{\text{format}}$ is given only when the model successfully follows the required format.
-This encourages the LLM to produce well-structured outputs, facilitating reliable parsing and consistent downstream evaluation during training.
 Answer Reward.
@@
 This combined reward jointly encourages the model to produce well-structured outputs, generate correct answers, and discover diverse and valid reasoning paths during reinforcement learning.
-We optimize RoE using Group Relative Policy Optimization (GRPO) (Guo et al., 2025b), a variant of Proximal Policy Optimization (PPO) (Schulman et al., 2017).
-GRPO eliminates the need for an additional value function approximation by using the average reward of multiple sampled outputs as a relative baseline.
-For each input question $x$, GRPO samples a group of outputs $\{y_{1},y_{2},\ldots,y_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$ and optimizes the new policy $\pi_{\theta}$ by maximizing the following objective:
-$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{x\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\min\Big(r_{i,t}(\theta)\,\hat{A}_{i,t},$
-$\quad\mathrm{clip}\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,\hat{A}_{i,t}\Big)-\gamma D_{\text{KL}}\!\big[\pi_{\theta}(\cdot|x)\,\|\,\pi_{\text{ref}}(\cdot|x)\big]\Bigg]$
-where
-$r_{i,t}(\theta)=\frac{\pi_{\theta}(y_{i,t}\mid x,y_{i,<t})}{\pi_{\theta_{\text{old}}}(y_{i,t}\mid x,y_{i,<t})},\qquad\hat{A}_{i,t}=\frac{R_{i}-\mathrm{mean}(\{R_{j}\}_{j=1}^{G})}{\mathrm{std}(\{R_{j}\}_{j=1}^{G})},$
-and $\gamma$ is a hyperparameter balancing task-specific rewards and KL-divergence regularization.
-This formulation encourages the model to prefer outputs with higher relative rewards within each group, stabilizing training while improving sample efficiency.
-In our case, the rewards $R_{i}$ correspond to the total rule-based rewards $\mathcal{R}_{\text{total}}$.
+We optimize RoE using Group Relative Policy Optimization (GRPO) (Guo et al., 2025a), a variant of Proximal Policy Optimization (PPO) (Schulman et al., 2017).
 In summary, RoE is first supervised through step-wise fine-tuning (Stage I) to learn the fundamental exploration behavior from gold trajectories, and is then refined via reinforcement learning (Stage II) to improve its reasoning and generalization ability.
@@
 To evaluate the effectiveness of RoE, we use two widely adopted datasets: WebQSP (Yih et al., 2016) and CWQ (Talmor and Berant, 2018), both constructed upon the Freebase knowledge graph (Bollacker et al., 2008).
-To further assess the generalization ability of GraphRAG methods across different knowledge graphs, we additionally use MetaQA (Zhang et al., 2018), which is built on the WikiMovies KG (Miller et al., 2016), and PathQuestion (Zhou et al., 2018), which uses a Freebase subset with modified relation names.
-For cross-graph generalization evaluation, we randomly sample 1,000 multi-hop questions from each of these two datasets.
+To further assess the generalization ability, we additionally use MetaQA (Zhang et al., 2018), which is built on the WikiMovies KG (Miller et al., 2016), and PathQuestion (Zhou et al., 2018), which uses a Freebase subset with modified relation names.
 Evaluation Metrics.
-We follow previous works (He et al., 2021; Guo et al., 2025a; Li et al., 2024) and adopt two evaluation metrics: Hit and F1.
+We follow previous works (He et al., 2021; Guo et al., 2025b; Li et al., 2024) and adopt two evaluation metrics: Hit and F1.
 Specifically, the Hit measures the proportion of questions for which at least one correct predicted answer, while F1 provides a balanced evaluation of both precision and recall.
@@
 (2) GNN-Only: models that use GNNs for both retrieval and reasoning.
-We include GraftNet (Sun et al., 2018), SR+NSM (Zhang et al., 2022b), and UniKGQA (Jiang et al., 2022).
+We include GraftNet (Sun et al., 2018), SR+NSM (Zhang et al., 2022a), and UniKGQA (Jiang et al., 2022).
 (3) LLMs + GNNs: hybrid approaches that combine LLMs and GNNs for retrieval and generation.
@@
 (4) LLMs + KG: methods that employ LLMs for both retrieval and generation over knowledge graphs.
-We compare with ToG (Sun et al., 2023), RoG (Luo et al., 2023), KD-CoT (Wang et al., 2023a), and SubGraphRAG (Li et al., 2024).
+We compare with ToG (Sun et al., 2023), RoG (Luo et al., 2023), KD-CoT (Wang et al., 2023b), and SubGraphRAG (Li et al., 2024).
 RoE settings.
@@
 Moreover, to eliminate the influence of backbone model, we use the same Llama-3.1-8B-Instruct model (Touvron et al., 2023) for most LLM-based baselines, including RoG, KD-CoT, SubGraphRAG, G-Retriever, and GNN-RAG.
+We also evaluate with different backbone models in Appendix D.
 5.2. Overall Performance Comparison
@@
 On the CWQ dataset, RoE also achieves a 3.1% relative improvement in Hit, demonstrating its strong ability to reason on graphs.
-  • The LLM-only models, which lack access to external graph information, generally perform worse than methods that incorporate graphs.
+  • The LLM-only models, which lack access to graph information, generally perform worse than methods that incorporate graphs.
 However, stronger models, such as GPT-4.1, achieve better results, suggesting that large LLMs already possess internal knowledge relevant to the benchmarks.
@@
 We train these variants on WebQSP dataset and test their performance on WebQSP, MetaQA and PathQuestion dataset.
-Figure 4. Performance of RoE and its variants pretrained on the WebQSP dataset. Marker shapes denote different model variants, while filled and hollow markers represent Hit and metrics, respectively.
+Figure 4.
+Performance of RoE and its variants pretrained on the WebQSP dataset.
+Marker shapes denote different model variants, while filled and hollow markers represent Hit and F1 metrics, respectively.
 The results are presented in Figure 4, from which we make the following observations: (1) The RL stage is essential for RoE, especially under transfer settings.
@@
 Extensive experiments on multiple KGQA benchmarks demonstrate that RoE consistently outperforms strong baselines and generalizes effectively across different question distributions and graphs.
+Acknowledgment
+This work is supported by the National Science Foundation (NSF) under grant numbers CNS2321416, IIS2212032, IIS2212144, IIS2504089, DUE2234015, CNS2246050, DRL2405483, IIS 2524379, and NAIRR 250188, the Michigan Department of Agriculture and Rural Development, US Dept of Commerce, Gates Foundation, Amazon Faculty Award, Meta, NVIDIA, Microsoft and SNAP.
 Appendix A Algorithm of SFT Dataset Construction
@@
 For reproducibility, all random seeds are fixed, and the same prompt templates for SFT and RL.
+Appendix D Results with different backbone models
+To evaluate robustness across model sizes, we additionally train RoE-3B and other baseline models with Llama-3.2-3B-Instruct backbone.
+All models are trained on WebQSP and evaluated on both WebQSP and MetaQA.
 Algorithm 1 SFT Dataset Construction
@@
 28:return $\mathcal{D}_{\text{SFT}}$
+Table 4. Performance comparison on the WebQSP and MetaQA datasets. All models use the Llama-3.2-3B-Instruct backbone and are trained on the WebQSP dataset.
+WebQSP $\rightarrow$ WebQSP WebQSP $\rightarrow$ MetaQA
+Method Hit F1 Hit F1
+G-Retrieval-3B 67.32 47.25 12.7 5.58
+RoG-3B 73.95 55.79 10.80 6.00
+GNN-RAG-3B 81.51 67.16 41.70 13.94
+RoE-3B 85.56 70.62 74.20 57.12
+From the results in Table 4, RoE continues to outperform all baselines by a large margin, especially in the transfer setting on MetaQA.
+Notably, RoE-3B surpasses baselines even when they use Llama-8B, especially on the transfer setting.
+GRPO vs.
+GSPO yields similar performance, showing that RoE is not sensitive to the RL algorithm choice.